### v3

# VideoPainter: Any-length Video Inpainting and Editing with Plug-and-Play Context Control

Abstract.
Video inpainting, crucial for the media industry, aims to restore corrupted content. However, current methods relying on limited pixel propagation or single-branch image inpainting architectures face challenges with generating fully masked objects, balancing background preservation with foreground generation, and maintaining ID consistency over long video. To address these issues, we propose VideoPainter, an efficient dual-branch framework featuring a lightweight context encoder. This plug-and-play encoder processes masked videos and injects background guidance into any pre-trained video diffusion transformer, generalizing across arbitrary mask types, enhancing background integration and foreground generation, and enabling user-customized control. We further introduce a strategy to resample inpainting regions for maintaining ID consistency in any-length video inpainting. Additionally, we develop a scalable dataset pipeline using advanced vision models and construct VPData and VPBench—the largest video inpainting dataset with segmentation masks and dense caption (¿390K clips) —to support large-scale training and evaluation. We also show VideoPainter’s promising potential in downstream applications such as video editing. Extensive experiments demonstrate VideoPainter’s state-of-the-art performance in any-length video inpainting and editing across $8$ key metrics, including video quality, mask region preservation, and textual coherence.
Artificial Intelligence Generative Content, Computer Vision, Video Inpainting, Video Editing
submissionid: 626journal: TOGccs: Computing methodologies Computer vision
1. Introduction
Video inpainting (Quan et al., 2024), which aims to restore the corrupted video while maintaining coherence, facilitates numerous applications, including try-on (Fang et al., 2024), film production (Polyak et al., 2024), and video editing (Sun et al., 2024). Recently, Diffusion Transformers (DiT) (Peebles and Xie, 2023; OpenAI, 2024) have shown promise in video generation, leading to the exploration of generative video inpainting (Zhang et al., 2024b; Zi et al., 2024).
Figure 2. Framework Comparison. Non-generative approaches, limited to pixel propagation from backgrounds, fail to inpaint fully segmentation-masked objects. Generative methods adapt single-branch image inpainting models to video by adding temporal attention, struggling to maintain background fidelity and generate foreground contents in one model. In contrast, VideoPainter implements a dual-branch architecture that leverages an efficient context encoder with any pre-trained DiT, decoupling video inpainting to background preservation and foreground generation, and enabling plug-and-play video inpainting control.
Existing approaches, as illustrated in Fig. 2, can be broadly categorized into two types: (1) Non-Generative methods (Zhou et al., 2023; Li et al., 2022; Lee et al., 2019) depend on limited pixel feature propagation (physical constraints or model architectural priors), which only take masked videos as inputs and cannot generate fully segmentation-masked objects. (2) Generative methods (Zhang et al., 2024b; Zi et al., 2024; Wang et al., 2024) extend single-branch image inpainting architectures (Rombach and Esser, 2022) to video by incorporating temporal attention, which struggles to balance background preservation and foreground generation in one model and obtain inferior temporal coherence compared to native video DiTs. Moreover, both paradigms neglect long video inpainting and struggle to maintain consistent object identity with long videos.
This motivates us to decompose video inpainting into background preservation and foreground generation and adopt a dual-branch architecture in DiTs, where we can incorporate a dedicated context encoder for masked video feature extraction while utilizing the pre-trained DiT’s capabilities to generate semantic coherent video content conditioned on both the preserved background and text prompts. Similar observations have been made in image inpainting research, notably in BrushNet (Ju et al., 2024) and ControlNet (Zhang et al., 2023). However, directly applying their architecture to video DiTs presents several challenges: (1) Given Video DiT’s robust generative foundation and heavy model size, replicating the full/half-giant Video DiT backbone as the context encoder would be unnecessary and computationally prohibitive. (2) Unlike BrushNet’s pure convolutional control branch, DiT’s tokens in masked regions inherently contain background information due to global attention, complicating the distinction between masked and unmasked regions in DiT backbones. (3) ControlNet lacks feature injection across all layers, hindering dense background control for inpainting tasks.
To address these challenges, we introduce VideoPainter, which enhances pre-trained DiT with a lightweight context encoder comprising only 6% of the backbone parameters, to form the first efficient dual-branch video inpainting architecture. VideoPainter features three main components: (1) A streamlined context encoder with just two layers, which integrates context features into the pre-trained DiT in a group-wise manner, ensuring efficient and dense background guidance. (2) Mask-selective feature integration to clearly distinguish the tokens of the masked and unmasked region. (3) A novel inpainting region ID resampling technique to efficiently process videos of any length while maintaining ID coherence. By freezing the pre-trained context encoder and DiT backbone, and adding an ID-Adapter, we enhance the backbone’s attention sampling by concatenating the original key-value vectors with the inpainting region tokens. During inference, inpainting region tokens from previous clips are appended to the current key-value vectors, ensuring the long-term preservation of target IDs. Notably, our VideoPainter supports plug-and-play and user-customized control.
For large-scale training, we develop a scalable dataset pipeline using advanced vision models (OpenAI, 2024; Ravi et al., 2024; Zhang et al., 2024a), constructing the largest video inpainting dataset, VPData, and benchmark, VPBench, with over 390K clips featuring precise segmentation masks and dense text captions. We further demonstrate VideoPainter’s potential by establishing an inpainting-based video editing pipeline that delivers promising results.
To validate our approach, we compare VideoPainter against previous state-of-the-art (SOTA) baselines and a single-branch fine-tuning setup that combines noisy latent, masked video latent, and mask at the input channel. VideoPainter demonstrates superior performance in both training efficiency and final results.
In summary, our contributions are as follows:
  • We propose VideoPainter, the first dual-branch video inpainting framework that supports plug-and-play background controls.
  • We design a lightweight context encoder for efficient and dense background control, and inpainting region ID resampling for ID consistency in any-length video inpainting and editing.
[…]
  • Experiments show VideoPainter achieves state-of-the-art performance across $8$ metrics including video quality, masked region preservation, and text alignment in video inpainting and editing.
2. Related Work
2.1. Video Inpainting
Video inpainting approaches can be broadly classified into two categories based on whether they possess generative capabilities:
Non-generative methods.
These methods (Zhou et al., 2023; Li et al., 2022; Hu et al., 2020; Zhang et al., 2022a, b) leverage architecture priors to facilitate pixel propagation. This includes utilizing local perception of 3D CNNs (Chang et al., 2019a; Wang et al., 2019; Hu et al., 2020; Chang et al., 2019b), and exploiting the global perception of attention to retrieve and aggregate tokens with similar texture for filling masked video (Lee et al., 2019; Zeng et al., 2020; Liu et al., 2021; Zhang et al., 2022a). They also introduce various physical quantities, especially optical flow, as auxiliary conditions as it simplifies RGB pixel inpainting by completing less complex flow fields (Kim et al., 2019; Li et al., 2020; Zou et al., 2021; Xu et al., 2019; Gao et al., 2020; Zhang et al., 2022c, b). However, they are only effective for partial object occlusions with random masks but face significant limitations when inpaint fully masked regions due to insufficient contexts.
Generative methods.
Recent advances in generative foundation models (Rombach et al., 2022; Guo et al., 2023) have sparked numerous approaches that leverage additional modules or training strategies to extend backbones’ capabilities for video inpainting (Zhang et al., 2024b; Zi et al., 2024; Wang et al., 2024). AVID (Zhang et al., 2024b) and COCOCO (Zi et al., 2024) represent the most related recent works. Both adopt a similar implementation by augmenting Stable Diffusion Inpainting (Rombach and Esser, 2022) with trainable temporal attention layers. This architecture includes per-frame region filling based on the image inpainting backbone and temporal smoothing with temporal attention. Despite showing promising results for both random and segmentation masks due to their generative abilities, they struggle to balance background preservation and foreground generation with text caption (Ju et al., 2024; Li et al., 2024) within the single backbone. AVID also explores any-length video inpainting by smoothing latent at segment boundaries and using the middle frame as the ID reference. In contrast, VideoPainter is a dual-branch framework by decoupling video inpainting into foreground generation and background-guided preservation. It employs an efficient context encoder to guide any pre-trained DiT, facilitating plug-and-play control. Furthermore, VideoPainter also introduces a novel inpainting region ID resampling technique that enables ID consistency in any-length video inpainting.
Table 1. Comparison of video inpainting datasets. Our VPData is the largest video inpainting dataset to date, comprising over 390K high-quality clips with segmentation masks, video captions, and masked region descriptions.
Dataset #Clips Duration Video Caption Masked Region Desc.
DAVIS (Perazzi et al., 2016) 0.4K 0.1h
YouTube-VOS (Xu et al., 2018) 4.5K 5.6h
VOST (Tokmakov et al., 2023) 1.5K 4.2h
MOSE (Ding et al., 2023) 5.2K 7.4h
LVOS (Hong et al., 2023) 1.0K 18.9h
SA-V (Ravi et al., 2024) 642.6K 196.0h
Ours 390.3K 866.7h
2.2. Video Inpainting Datasets
Recent advances in segmentation (Ravi et al., 2024) have created many video segmentation datasets (Perazzi et al., 2016; Xu et al., 2018; Hong et al., 2023; Ding et al., 2023; Tokmakov et al., 2023; Darkhalil et al., 2022). Among these, DAVIS (Perazzi et al., 2016) and YouTube-VOS (Xu et al., 2018) have become prominent benchmarks for video inpainting due to their high-quality masks and diverse object categories. However, the existing datasets face two primary limitations: (1) insufficient scale to meet the data requirements of generative models, and (2) the absence of crucial control conditions necessary for generating masked objects such as video captions. In contrast, as shown in Tab. 10, we developed a scalable dataset pipeline based on state-of-the-art vision understanding models (Zhang et al., 2024a; Ravi et al., 2024; OpenAI, 2024), and constructed the largest video inpainting dataset to date with over 390K clips, each annotated with segmentation masks and dense video captions.
Figure 3. Dataset Construction Pipeline. It consists of five pre-processing steps: collection, annotation, splitting, selection, and captioning.
3. Method
Sec. 3.1 and Fig. 3 illustrate our pipeline for building VPData and VPBench. Sec. 3.2 and Fig. 4 show our dual-branch VideoPainter. Sec. 3.3 and Sec. 3.4 introduce our inpainting region ID resampling approach for any-length video inpainting and plug-and-play control.
Figure 4. Model overview. The upper figure shows the architecture of VideoPainter. The context encoder performs video inpainting based on concatenation of the noisy latent, downsampled masks, and masked video latent via VAE. Features extracted by the context encoder are integrated into the pre-trained DiT in a group-wise and token-selective manner, where two encoder layers modulate the first and second halves of the DiT, respectively, and only the background tokens will be integrated into the backbone to prevent information ambiguity. The lower figure illustrates the inpainting ID region resampling with the ID Resample Adapter. During training, tokens of the current masked region are concatenated to the KV vectors, enhancing ID preservation of the inpainting region. During inference, the ID tokens of the last clip are concatenated to the current KV vectors, maintaining ID consistency with the last clip by resampling.
3.1. VPData and VPBench Construction Pipeline
To address the challenges of limited size and lack of text annotations, we present a scalable dataset pipeline leveraging advanced vision models (Ravi et al., 2024; OpenAI, 2024; Zhang et al., 2024a). This leads to VPData and VPBench, the largest video inpainting dataset and benchmark with precise masks and video/masked region captions. As shown in Fig. 3, the pipeline involves $5$ steps: collection, annotation, splitting, selection, and captioning.
Collection. We chose Videvo and Pexels  [^1] Videvo: https://www.videvo.net/, Pexels: https://www.pexels.com/ as our data sources. We finally obtained around $450K$ videos from these sources.
Annotation. For each collected video, we implement a cascaded workflow for automated annotation:
  • We employ the Recognize Anything Model (Zhang et al., 2024a) for open-set video tagging to identify primary objects.
  • Based on the detected object tags, we utilize Grounding DINO (Liu et al., 2023) to detect bounding boxes for objects at fixed intervals.
  • These bounding boxes serve as prompts for SAM2 (Ravi et al., 2024), which generates high-quality mask segmentations.
  • Then we employ rigorous filtering criteria: inter-frame mask area variation $\Delta<20\%$ and frame coverage maintained between $30\%-70\%$ to ensure reliable segmentation masks quality.
Splitting. Scene transitions may occur while tracking the same object from different angles, causing disruptive view changes. We utilize PySceneDetect (Castellano, 2024) to identify scene transitions and subsequently partition the masks. Then we segmented the sequences into 10-second intervals and discarded short clips (¡ 6s).
Selection. We employ $3$ key criteria: (1) Aesthetic Quality, evaluated using the Laion-Aesthetic Score Predictor (Schuhmann et al., 2022); (2) Motion Strength, predicted by optical flow measurements using the RAFT(Teed and Deng, 2020); and (3) Content Safety, assessed via the Stable Diffusion Safety Checker (Rombach et al., 2022).
Captioning. As Tab. 10 shows, existing video segmentation datasets lack textual annotations, primary conditions in generation (Chen et al., 2023; Betker et al., 2023), creating a data bottleneck for applying generative models to video inpainting. Therefore, we leverage SOTA VLMs, specifically CogVLM2 (Wang et al., 2023) and GPT-4o (OpenAI, 2024), to uniformly sample keyframes and generate dense video captions and detailed descriptions of the masked objects.
3.2. Dual-branch Inpainting Control
We incorporate masked video features into the pre-trained diffusion transformer (DiT) via an efficient context encoder, to decouple the background context extraction and foreground generation. This encoder processes a concatenated input of noisy latent, masked video latent, and downsampled masks. Specifically, the noisy latent provides information about the current generation. The masked video latent, extracted via VAE, aligns with the pre-trained DiT’s latent distribution. We apply cubic interpolation to downsample masks, ensuring dimensional compatibility between masks and latents.
Based on DiT’s inherent generative abilities (OpenAI, 2024), the control branch only needs to extract contextual cues to guide the backbone in preserving background and generating foreground. Therefore, instead of previous heavy approaches that duplicate half or all of the backbone (Ju et al., 2024; Zhang et al., 2023), VideoPainter employs a lightweight design by cloning only the first two layers of pre-trained DiT, accounting for merely 6% of the backbone parameters. The pre-trained DiT weights provide a robust prior for extracting masked video features. The context encoder features are integrated into the frozen DiT in a group-wise, token-selective manner. The group-wise feature integration is formulated as follows: the first layer’s features are added back to the initial half of the backbone, while the second layer’s features are integrated into the latter half, achieving lightweight and efficient context control. The token-selective mechanism is a pre-filtering process, where only tokens representing pure background are added back, while others are excluded from integration, as shown in the upper right of Fig. 4. This ensures that only the background context is fused into the backbone, preventing potential ambiguity during backbone generation.
[…]
$\epsilon_{\theta}\left(z_{t},t,C\right)_{i}=\epsilon_{\theta}\left(z_{t},t,C\right)_{i}+\mathcal{Z}\left(\epsilon_{\theta}^{VideoPainter}\left(\left[z_{t},z_{0}^{masked},m^{resized}\right],t\right)_{i//\frac{n}{2}}\right)$
3.3. Target Region ID Resampling
While current DiTs show promise in handling temporal dynamics (Kuaishou, 2024; Bian et al., 2024a), they struggle to maintain smooth transitions and long-term identity consistency.
Smooth Transition. Following AVID (Zhang et al., 2024b), we employ overlapping generation and weighted average to maintain consistent transitions. Additionally, we utilize the last frame of the previous clip (before overlap) as the first frame of the current clip’s overlapping region to ensure visual appearance continuity.
Identity Consistency. To maintain identity consistency in the long video, we introduce an inpainting region ID resampling method, as shown in lower Fig. 4. During training, we freeze both the DiT and the context encoder. Then we add trainable ID-Resample Adapters into the frozen DiT (LoRA), enabling ID resampling functionality. Specifically, tokens from the current masked region, which contain the desired ID, are concatenated with the KV vectors, thereby enhancing ID preservation in the inpainting region through additional KV resampling. During inference, we prioritize maintaining ID consistency with the inpainting region tokens from the previous clip, as it represents the most temporally proximate generated result. Specifically, given current $Q^{v}_{i}$, $K^{v}_{i}$, and $V^{v}_{i}$, we concatenate tokens containing ID information ($K^{id}_{i}$ and $V^{id}_{i}$) to the current KV pairs (during training, these are tokens from the current inpainting region; during inference, from the previous clip’s inpainting region). This forms new KV-vectors $[K^{v}_{i},K^{id}_{i}]$ and $[V^{v}_{i},V^{id}_{i}]$ (where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes concatenation), enabling the model to sample necessary ID information and better maintain ID consistency.
3.4. Plug-and-Play Control
Our plug-and-play framework demonstrates versatility across two aspects: it supports various stylization backbones or LoRAs and is compatible with both text-to-video (T2V) (Yang et al., 2024; NVIDIA, 2025) and image-to-video (I2V) (Guo et al., 2024; Shi et al., 2024) DiT architectures. The I2V compatibility particularly enables seamless integration with existing image inpainting capabilities. When utilizing an I2V DiT backbone, VideoPainter requires only one additional step: generating the initial frame using any image inpainting model guided by the masked region’s text caption. This inpainted frame then serves as both the image condition and the first masked video frame. These capabilities further demonstrate the exceptional transferability and versatility of VideoPainter.
4. Experiments
4.1. Implementation details
VideoPainter is built upon a pre-trained Image-to-Video Diffusion Transformer CogVideo-5B-I2V (Yang et al., 2024) (by default) and its Text-to-Video version. In training, we use VPData at a $480\times 720$ resolution, learning rate $1\times 10^{-5}$, batch size $1$ for both the context encoder ($80,000$ steps) and the ID Resample Adapter ($2,000$ steps) in two stages with AdamW. In training, we randomly sample dilation and erosion with kernel sizes $\in[8,32]$ to enhance robustness to mask precision. This also enables our random-mask inpainting.
Benchmarks.
In video inpainting, we employ Davis (Perazzi et al., 2016) as the benchmark for random masks and VPBench for segmentation-based masks. VPBench consists of $100$ 6-second videos for standard video inpainting, and $16$ videos with an average duration of more than $30$ seconds for long video inpainting. The VPBench includes diverse content including objects, humans, animals, landscapes, and multi-range masks. For video editing evaluation, we also utilize VPBench, which includes four fundamental editing operations (add, remove, swap, and change) and comprises $45$ 6-second videos and $9$ videos with an average duration of $30$ seconds.
Metrics.
We consider $8$ metrics from three aspects: masked region preservation, text alignment, and video generation quality.
  • $\bullet$
    Masked Region Preservation. We follow previous works using standard PSNR (Wikipedia contributors, 2024c), LPIPS (Zhang et al., 2018), SSIM (Wang et al., 2004), MSE (Wikipedia contributors, 2024b) and MAE (Wikipedia contributors, 2024a) in the unmasked region among the generated video and the original video.
  • $\bullet$
    Text Alignment. We employ CLIP Similarity (CLIP Sim) (Wu et al., 2021) to assess the semantic consistency between the generated video and its corresponding text caption. We also measure CLIP Similarity within the masked regions (CLIP Sim (M)).
  • $\bullet$
    Video Generation Quality. Following previous methods, we use FVID (Wang et al., 2018) to measure the generated video quality.
4.2. Video Inpainting
Figure 5. Comparison of previous inpainting methods and VideoPainter on standard and long video inpainting. More visualizations are in the demo video.
Quantitative comparisons
Tab. 2 shows the quantitative comparison on VPBench and Davis (Perazzi et al., 2016). We compare the inpainting results of non-generative ProPainter (Zhou et al., 2023), generative COCOCO (Zi et al., 2024), and Cog-Inp (Yang et al., 2024), a strong baseline proposed by us, which inpaint first frame using image inpainting models and use the I2V backbone to propagate results with the latent blending operation (Avrahami et al., 2023). In the segmentation-based VPBench, ProPainter, and COCOCO exhibit the worst performance across most metrics, primarily due to the inability to inpaint fully masked objects and the single-backbone architecture’s difficulty in balancing the competing background preservation and foreground generation, respectively. In the random mask benchmark Davis, ProPainter shows improvement by leveraging partial background information. However, VideoPainter achieves optimal performance across segmentation (standard and long length) and random masks through its dual-branch architecture that effectively decouples background preservation and foreground generation.
Qualitative comparisons
The qualitative comparison with previous video inpainting methods is shown in Fig. 5. VideoPainter consistently shows exceptional results in video coherence, quality, and alignment with text caption. Notably, ProPainter fails to generate fully masked objects because it only depends on background pixel propagation instead of generating. While COCOCO demonstrates basic functionality, it fails to maintain consistent ID in inpainted regions ( inconsistent vessel appearances and abrupt terrain changes) due to its single-backbone architecture attempting to balance background preservation and foreground generation. Cog-Inp achieves basic inpainting results; however, its blending operation’s inability to detect mask boundaries leads to significant artifacts. Moreover, VideoPainter can generate coherent videos exceeding one minute while maintaining ID consistency through our ID resampling.
Table 2. Quantitative comparisons among VideoPainter and other video inpainting models in VPBench for segmentation mask (Standard (S) and Long (L) Video) and Davis for random mask: ProPainter (Zhou et al., 2023), COCOCO (Zi et al., 2024), and Cog-Inp (Yang et al., 2024). Metrics include masked region preservation, text alignment, and video quality. Red stands for the best, Blue stands for the second best.
Metrics Masked Region Preservation Text Alignment Video Quality
[…]
VPBench-S ProPainter 20.97 0.87 9.89 1.24 3.56 7.31 17.18 0.44
COCOCO 19.27 0.67 14.80 1.62 6.38 7.95 20.03 0.69
Cog-Inp 22.15 0.82 9.56 0.88 3.92 8.41 21.27 0.18
Ours 23.32 0.89 6.85 0.82 2.62 8.66 21.49 0.15
VPBench-L ProPainter 20.11 0.84 11.18 1.17 3.71 9.44 17.68 0.48
COCOCO 19.51 0.66 16.17 1.29 6.02 11.00 20.42 0.62
Cog-Inp 19.78 0.73 12.53 1.33 5.13 11.47 21.22 0.21
Ours 22.19 0.85 9.14 0.71 2.92 11.52 21.54 0.17
Davis ProPainter 23.99 0.92 5.86 0.98 2.48 7.54 16.69 0.12
COCOCO 21.34 0.66 10.51 0.92 4.99 6.73 17.50 0.33
Cog-Inp 23.92 0.79 10.78 0.47 3.23 7.03 17.53 0.17
Ours 25.27 0.94 4.29 0.45 1.41 7.21 18.46 0.09
4.3. Video Editing
VideoPainter can be used for video inpainting by employing Vison Language Models (OpenAI, 2024; Team et al., 2024) to generate modified captions based on user editing instructions and source captions and apply VideoPainter to inpaint based on the modified captions. Tab. 3 shows the quantitative comparison on VPBench. We compare the editing results of inverse-based UniEdit (Bai et al., 2024), DiT-based DiTCtrl (Cai et al., 2024), and end-to-end ReVideo (Mou et al., 2024). For both standard and long videos in VPBench, VideoPainter achieves superior performance, even surpassing the end-to-end ReVideo. This success can be attributed to its dual-branch architecture, which ensures excellent background preservation and foreground generation capabilities, maintaining high fidelity in non-edited regions while ensuring edited regions closely align with editing instructions, complemented by inpainting region ID resampling that maintains ID consistency in long video. The qualitative comparison with previous video inpainting methods is shown in Fig. 5. VideoPainter demonstrates superior performance in preserving visual fidelity and text-prompt consistency.
Table 3. Quantitative comparisons among VideoPainter and other video editing models in VPBench (Standard and Long Video): UniEdit (Bai et al., 2024), DitCtrl (Cai et al., 2024), and ReVideo (Mou et al., 2024). Metrics include masked region preservation, text alignment, and video quality. Red stands for the best, Blue stands for the second best.
Metrics Masked Region Preservation Text Alignment Video Quality
[…]
Standard UniEdit 9.96 0.36 56.68 11.08 25.78 8.46 14.23 1.36
DitCtrl 9.30 0.33 57.42 12.73 27.45 8.52 15.59 0.57
ReVideo 15.52 0.49 27.68 3.49 11.14 9.34 20.01 0.42
Ours 22.63 0.91 7.65 1.02 2.90 8.67 20.20 0.18
Long UniEdit 10.37 0.30 54.61 10.25 24.89 10.85 15.42 1.00
DitCtrl 9.76 0.28 62.49 11.50 26.64 11.78 16.52 0.56
ReVideo 15.50 0.46 28.57 3.92 12.24 11.22 20.50 0.35
Ours 22.60 0.90 7.53 0.86 2.76 11.85 19.38 0.11
Figure 6. Comparison of previous editing methods and VideoPainter on standard and long video editing. More visualizations are in the demo video.
Table 4. User Study: User preference ratios comparing VideoPainter with video inpainting and editing baselines. For each sample, participants selected only one model that produced the best results for each criterion. We evaluate performance using the average proportion of being selected as the best response. For video inpainting, we compared VideoPainter against ProPainter (Zhou et al., 2023), COCOCO (Zi et al., 2024), and Cog-Inp (Yang et al., 2024). For video editing, we compared VideoPainter against UniEdit (Bai et al., 2024), DitCtrl (Cai et al., 2024), and ReVideo (Mou et al., 2024). Detailed results are in the appendix.
Task Video Inpainting Video Editing
Background Text Video Background Text Video
Preservation Alignment Quality Preservation Alignment Quality
Ours $74.2\%$ $82.5\%$ $87.4\%$ $78.4\%$ $76.1\%$ $81.7\%$
4.4. Human Evaluation
We conducted a user study on video inpainting and editing tasks using standard-length video samples from the VPBench inpainting and editing subsets. Thirty participants evaluated 50 randomly selected cases based on background preservation, text alignment, and video quality. As shown in Tab. 4, VideoPainter significantly outperformed existing baselines, achieving higher preference rates across all evaluation criteria in both tasks. Detailed experiment settings and results are provided in the Appendix.
4.5. Ablation Analysis
We ablate on VideoPainter in Tab .7, including architecture, context encoder size, control strategy, and inpainting region ID resampling.
Table 5. Ablation Studies on VPBench. Single-Branch: We add input channels to adapt masked video and finetune the backbone. Layer Configuration (VideoPainter (*)): We vary the context encoder depth from one to four layers. w/o Selective Token Integration (w/o Select):: We bypass the token pre-selection step and integrate all context encoder tokens into DiT. T2V Backbone (VideoPainter (T2V)): We replace the backbone from image-to-video DiTs to text-to-video DiTs. w/o target region ID resampling (w/o Resample): We ablate on the target region ID resampling. (L) denotes evaluation on the long video subset. Red stands for the best result.
Metrics Masked Region Preservation Text Alignment Video Quality
[…]
Single-Branch 20.54 0.79 10.48 0.94 4.16 8.19 19.31 0.22
VideoPainter (1) 21.92 0.81 8.78 0.89 3.26 8.44 20.79 0.17
VideoPainter (4) 22.86 0.85 6.51 0.83 2.86 9.12 20.49 0.16
w/o Select 20.94 0.74 7.90 0.95 3.87 8.26 17.84 0.25
VideoPainter (T2V) 23.01 0.87 6.94 0.89 2.65 9.41 20.66 0.16
VideoPainter 23.32 0.89 6.85 0.82 2.62 8.66 21.49 0.15
w/o Resample (L) 21.79 0.84 8.65 0.81 3.10 11.35 20.68 0.19
VideoPainter (L) 22.19 0.85 9.14 0.71 2.92 11.52 21.54 0.17
Based on rows 1 and 5, the dual-branch VideoPainter significantly outperforms its single-branch counterpart by explicitly decoupling background preservation from foreground generation, thereby reducing model complexity and avoiding the trade-off between competing objectives in a single branch. Row 2 to row 6 of Tab. 7 demonstrate the rationale behind our key design choices: 1 utilizing a two-layer structure as an optimal balance between performance and efficiency for the context encoder, 2 implementing token-selective feature fusion based on segmentation mask information to prevent confusion from indistinguishable foreground-background tokens in the backbone, and 3 adapting plug-and-play control to different backbones with comparable performance. Furthermore, rows 7 and 8 verify the importance of employing inpainting region ID resampling for long videos, which maintains ID consistency by explicitly resampling inpainted region tokens from previous clips.
4.6. Plug-and-Play Control Ability
Fig. 7 demonstrates the flexible plug-and-play control capabilities of VideoPainter in base diffusion transformer selection. We showcase how VideoPainter can be seamlessly integrated with community-developed Gromit-style LoRA. Despite the significant domain gap between anime-style data and our training dataset, VideoPainter’s dual-branch architecture ensures its plug-and-play inpainting abilities, enabling users to select the most appropriate base model for specific inpainting requirements and expected results.
Figure 7. Integrating VideoPainter to Gromit-style LoRA (Cseti, 2024).
5. Discussion
In this paper, we introduce VideoPainter, the first dual-branch video inpainting framework with plug-and-play control capabilities. Our approach features three key innovations: (1) a lightweight plug-and-play context encoder compatible with any pre-trained video DiTs, (2) an inpainting region ID resampling technique for maintaining long video ID consistency, and (3) a scalable dataset pipeline that produced VPData and VPBench, containing over 390K video clips with precise masks and dense captions. VideoPainter also shows promise in video editing applications. Extensive experiments demonstrate that VideoPainter achieves state-of-the-art performance across 8 metrics in video inpainting and editing, particularly in video quality, mask region preservation, and text coherence.
However, VideoPainter still has limitations: (1) Generation quality is limited by the base model, which may struggle with complex physical and motion modeling, and (2) performance is suboptimal with low-quality masks or misaligned video captions.
Acknowledgements
This work was supported in part by the CUHK Strategic Seed Funding for Collaborative Research Scheme under Grant No. 3136023 and in part by the CUHK Research Matching Scheme under Grant No. 7106937, 8601130, and 8601440.
Appendix A User Study
For a comprehensive evaluation beyond the quantitative and qualitative experiments presented in sections 4.2 and 4.3, we conducted a user study on both video inpainting and editing tasks using standard-length video samples from the VPBench inpainting subset and editing subset. The study involved 30 participants evaluating 50 randomly selected cases across three criteria: background preservation, text alignment, and video quality. Specifically, we assessed (1) Background Preservation: the error distance between the uninpainted regions of the generated video and the original video, (2) Text Alignment: the semantic coherence between inpainted regions and text captions, and (3) Video Quality: the overall visual fidelity of the generated results. As shown in Tab. 6, our approach demonstrates superior performance across all three aspects in both tasks. For video inpainting, our method achieves win rates of 74.2%, 82.5%, and 87.4% respectively. Similarly, for video editing, we obtain win rates of 78.4%, 76.1%, and 81.7% across the three aspects.
Table 6. User study evaluation comparing VideoPainter against state-of-the-art video inpainting and editing models. We conducted comprehensive comparisons on the VPBench, randomly sampling 50 examples from each of the inpainting and editing subsets. Human evaluators assessed the models’ outputs based on three criteria: background preservation, text alignment, and overall video quality. For each sample, participants selected only one model that produced the best results for each criterion. We evaluate performance using the proportion of model-generated outputs selected as the optimal response across all samples. For video inpainting, we compared against ProPainter (Zhou et al., 2023), COCOCO (Zi et al., 2024), and Cog-Inp (Yang et al., 2024). For video editing, we compared against UniEdit (Bai et al., 2024), DitCtrl (Cai et al., 2024), and ReVideo (Mou et al., 2024). Red stands for the best, Blue stands for the second best.
Metrics Background Text Video
[…]
Inpainting ProPainter 3.5 2.5 1.7
COCOCO 7.0 2.7 2.1
Cog-Inp 15.3 12.3 8.8
Ours 74.2 82.5 87.4
Editing UniEdit 3.3 3.7 2.8
DitCtrl 5.8 6.1 4.2
ReVideo 12.5 14.1 11.3
Ours 78.4 76.1 81.7
Appendix B Ablation on Dual-Branch Artecture
Table 7. Ablation Studies on VPBench. Single-Branch: We add input channels to adapt masked video and finetune the backbone. Layer Configuration (VideoPainter (*)): We vary the context encoder depth from one to four layers. w/o Selective Token Integration (w/o Select):: We bypass the token pre-selection step and integrate all context encoder tokens into DiT. T2V Backbone (VideoPainter (T2V)): We replace the backbone from image-to-video DiTs to text-to-video DiTs. w/o target region ID resampling (w/o Resample): We ablate on the target region ID resampling. (L) denotes evaluation on the long video subset. Red stands for the best result.
Metrics Masked Region Preservation Text Alignment Video Quality
[…]
Single-Branch 20.54 0.79 10.48 0.94 4.16 8.19 19.31 0.22
VideoPainter (1) 21.92 0.81 8.78 0.89 3.26 8.44 20.79 0.17
VideoPainter (4) 22.86 0.85 6.51 0.83 2.86 9.12 20.49 0.16
w/o Select 20.94 0.74 7.90 0.95 3.87 8.26 17.84 0.25
VideoPainter (T2V) 23.01 0.87 6.94 0.89 2.65 9.41 20.66 0.16
VideoPainter 23.32 0.89 6.85 0.82 2.62 8.66 21.49 0.15
w/o Resample (L) 21.79 0.84 8.65 0.81 3.10 11.35 20.68 0.19
VideoPainter (L) 22.19 0.85 9.14 0.71 2.92 11.52 21.54 0.17
Figure 8. Training loss curve of ablation of single branch fine-tuning and default dual branch VideoPainter. The training loss curves demonstrate that our dual-branch VideoPainter achieves superior convergence speed, stability, and final performance compared to single-branch fine-tuning, despite having significantly fewer trainable parameters.
As shown in Tab. LABEL:tab:ablation_on_branch and Fig. 8, we ablate on VideoPainter with a single branch fine-tuning setup.
Tab. LABEL:tab:ablation_on_branch clearly shows that the dual-branch VideoPainter significantly outperforms its single-branch counterpart by explicitly decoupling background preservation from foreground generation, thereby reducing model complexity and avoiding the trade-off between competing objectives in a single branch. The training loss curves in Fig. 8 demonstrate that VideoPainter, through effective dual-branch decoupling, achieves an advantageous optimization starting point and maintains stable, efficient convergence. By separating the competing objectives of background preservation and foreground generation via follow caption into distinct branches rather than a single backbone, VideoPainter achieves superior convergence performance, with a final loss approximately half that of single-branch finetuning.
Appendix C Ablation on Mask Quality
Table 8. Ablation Studies on VPBench: Kernel (*): We randomly sample dilation and erosion with varying kernel sizes $\in(8,16,32)$ for the segmentation masks. Kernel (Square): We randomly sample square masks with varying sizes $\in[8,32]$ and random locations. This reflects VideoPainter’s robustness to different mask qualities. Red stands for the best result.
Metrics Masked Region Preservation Text Alignment Video Quality
[…]
Kernel (8) 23.24 0.84 7.43 0.84 2.69 8.51 21.18 0.15
Kernel (16) 23.17 0.82 7.55 0.87 2.93 8.49 20.37 0.17
Kernel (32) 23.02 0.81 7.64 0.92 3.07 8.28 20.38 0.16
Kernel (Square) 22.46 0.73 9.37 0.94 3.75 7.81 19.88 0.21
VideoPainter 23.32 0.89 6.85 0.82 2.62 8.66 21.49 0.15
As shown in Tab. 8, we perform ablation studies on VideoPainter using a single branch fine-tuning configuration.
Tab. 8 clearly demonstrates that VideoPainter exhibits robust generalization across various qualities of video segmentation masks (which we augmented by randomly sampling different kernel sizes for dilation and erosion operations applied to the original segmentation masks). However, when utilizing random rectangular masks (which are highly atypical in video segmentation), model performance decreased by approximately 13%. This performance degradation primarily stems from artifacts generated along the rectangular edges, which we attribute to the substantial geometric disparity between these artificial rectangular masks and natural segmentation masks encountered in real-world scenarios.
Appendix D Ablation on ID Resampling
Table 9. Ablation Studies on VPBench: w/o target region ID resampling (w/o Resample): We ablate on the target region ID resampling. (L) denotes evaluation on the long video subset. Red stands for the best result.
Metrics Masked Region Preservation Text Alignment Video Quality
[…]
w/o Resample (L) 21.79 0.84 8.65 0.81 3.10 11.35 20.68 0.19
VideoPainter (L) 22.19 0.85 9.14 0.71 2.92 11.52 21.54 0.17
Figure 9. Long video inpainting comparison of original VideoPainter and w/o ID resampling ablation version.
As shown in Tab. 9 and Fig. 9, we ablate on VideoPainter with a w/o ID resampling ablation version.
Tab. LABEL:tab:ablation_on_branch clearly verifies the importance of employing inpainting region ID resampling for long videos, which maintains ID consistency by explicitly resampling inpainted region tokens from previous clips. Fig. 8 demonstrates that removing the ID resample adapter degrades long-video inpainting performance, with more severe deterioration as video length increases.
Appendix E More Related Works
E.1. Video Inpainting
Video inpainting approaches can be broadly classified into two categories based on whether they possess generative capabilities:
Non-generative methods.
These methods (Zhou et al., 2023; Li et al., 2022; Hu et al., 2020; Zhang et al., 2022a, b) leverage architecture priors to facilitate pixel propagation. This includes utilizing local perception of 3D CNNs (Chang et al., 2019a; Wang et al., 2019; Hu et al., 2020; Chang et al., 2019b), and exploiting the global perception of attention to retrieve and aggregate tokens with similar texture for filling masked video (Lee et al., 2019; Zeng et al., 2020; Liu et al., 2021; Zhang et al., 2022a). They also introduce various physical quantities, especially optical flow, as auxiliary conditions as it simplifies RGB pixel inpainting by completing less complex flow fields (Kim et al., 2019; Li et al., 2020; Zou et al., 2021; Xu et al., 2019; Gao et al., 2020; Zhang et al., 2022c, b). A notable exemplar of this approach is ProPainter (Zhou et al., 2023), which integrates flow completion models with spatiotemporal transformers into an end-to-end framework, achieving more faithful propagation through combined pixel and feature propagation with flow consistency verification. However, they are only effective for partial object occlusions with random masks but face significant limitations when inpaint fully masked regions due to insufficient contexts.
Generative methods.
Recent advances in generative foundation models (Rombach et al., 2022; Guo et al., 2023) have sparked numerous approaches that leverage additional modules or training strategies to extend backbones’ capabilities for video inpainting (Zhang et al., 2024b; Zi et al., 2024; Wang et al., 2024). VideoComposer (Wang et al., 2024), built upon a pre-trained video generation backbone Stable Video Diffusion (Blattmann et al., 2023), integrates various control signals (text, depth, mask, motion vectors) through a shared spatiotemporal condition fusion module. While it enables limited inpainting capabilities through mask and text conditioning, its performance in background preservation is compromised due to control condition compression and lack of masked video conditioning support. AVID (Zhang et al., 2024b) and COCOCO (Zi et al., 2024) represent the most related recent works. Both adopt a similar implementation by augmenting Stable Diffusion Inpainting (Rombach and Esser, 2022) with trainable temporal attention layers. This architecture includes per-frame region filling based on the image inpainting backbone and temporal smoothing with temporal attention. Despite showing promising results for both random and segmentation masks due to their generative abilities, they struggle to balance background preservation and foreground generation with text caption (Ju et al., 2024; Li et al., 2024; Bian et al., 2024b) within the single backbone. AVID also explores any-length video inpainting by smoothing latent at segment boundaries and using the middle frame as the ID reference. However, identity discontinuities occur during significant variations, and identity degradation emerges in longer video generation scenarios. In contrast, VideoPainter is a dual-branch framework by decoupling video inpainting into foreground generation and background-guided preservation. It employs an efficient context encoder to guide any pre-trained DiT, facilitating plug-and-play control and zero-shot adaptation across various stylization backbones—a capability absent in both non-generative and generative methods. Furthermore, VideoPainter also introduces a novel target region ID resampling technique that enables ID consistency in any-length video inpainting by explicitly resampling the inpainted region of the last generated clip.
Table 10. Comparison of video inpainting datasets. Our VPData is the largest video inpainting dataset to date, comprising over 390K high-quality clips with segmentation masks, video captions, and masked region descriptions.
Dataset #Clips Duration Video Caption Masked Region Desc.
DAVIS (Perazzi et al., 2016) 0.4K 0.1h
YouTube-VOS (Xu et al., 2018) 4.5K 5.6h
VOST (Tokmakov et al., 2023) 1.5K 4.2h
MOSE (Ding et al., 2023) 5.2K 7.4h
LVOS (Hong et al., 2023) 1.0K 18.9h
SA-V (Ravi et al., 2024) 642.6K 196.0h
Ours 390.3K 866.7h
E.2. Video Inpainting Datasets
Recent advances in segmentation (Ravi et al., 2024) have created many video segmentation datasets (Perazzi et al., 2016; Xu et al., 2018; Hong et al., 2023; Ding et al., 2023; Tokmakov et al., 2023; Darkhalil et al., 2022). Among these, DAVIS (Perazzi et al., 2016) and YouTube-VOS (Xu et al., 2018) have become prominent benchmarks for video inpainting due to their high-quality masks and diverse object categories. However, the existing datasets face two primary limitations: (1) insufficient scale to meet the data requirements of generative models, and (2) the absence of crucial control conditions necessary for generating masked objects such as video captions. These constraints have impeded the evaluation of state-of-the-art generative approaches, with methods like COCOCO and AVID resorting to random sampling from annotated internet video datasets for assessment. Consequently, these limitations have also hindered the further advancement of generative approaches in video inpainting. In contrast, as shown in Tab. 10, we developed a scalable dataset pipeline based on state-of-the-art vision understanding models (Zhang et al., 2024a; Ravi et al., 2024; OpenAI, 2024), and constructed the largest video inpainting dataset to date with over 390K clips, each annotated with segmentation masks and dense video captions. In contrast, we developed a scalable dataset construction pipeline based on state-of-the-art vision understanding models (Zhang et al., 2024a; Ravi et al., 2024; OpenAI, 2024), including open-set object detection, object segmentation mask tracking, shot transition detection, motion score and aesthetic quality filtering, and video annotation. By leveraging publicly available internet videos, we constructed the largest video inpainting dataset to date, comprising over 390K high-quality clips, each accompanied by precise segmentation masks, comprehensive video captions, and masked region descriptions, as Tab. 10 shows.
Appendix F More Discussions
As mentioned in the main paper, VideoPainter’s performance is suboptimal with low-quality or misaligned video captions. However, it should be noted that VideoPainter’s text-following capability (including complex physical or motion modeling abilities) derives from pre-trained DiTs. This can be enhanced by substituting more powerful DiT models, as our method is a plug-and-play framework independent of the internal DiT backbone architecture.
Appendix G More Visualizations
Figure 10. More video inpainting results.
Figure 11. More video editing results.